OpenStudy (anonymous):

A tent in the shape of a triangular prism has a square base with a side of 8 feet and a height of 6 feet. What is the volume of the tent?

OpenStudy (anonymous):

volume of pyramide = 1/3 *b*h = 1/3* 64*6=128